sreee:

Triangle XYZ was dilated by a scale factor of 2 to create triangle ACB and cos ∠X = 2 and 5 tenths over 5 and 59 hundredths.

Vocaloid:

Part A: Use complete sentences to explain the special relationship between the trigonometric ratios of triangles XYZ and ACB. You must show all work and calculations to receive full credit. (5 points) Part B: Explain how to find the measures of segments AC and AB. You must show all work and calculations to receive full credit. (5 points)

surjithayer:

\[\left| cosx \right|\le1\]

Vocaloid:

Keep in mind it says cosX 0 = (2 and 5 tenths) / (5 and 59 hundredths) not just 2

Sansthecomic:

calculate the sum I think then subtract...Maybe

surjithayer:

\[\cos x=\frac{ 2.5 }{5.59 }\] only thing we can say is \[\cos~A=\frac{ 2.5 }{ 5.59}\]